During take-off, an aircraft accelerates at 3.4 m/s2. The total mass of the aircraft is 8.0 × 103 kg. Calculate:

a) the net force acting on the aircraft during the acceleration.

b) the speed of the aircraft after 5.0 seconds of acceleration.


1
Expert's answer
2021-12-06T09:42:32-0500

(a)

"F_{net}=ma,""F_{net}=8.0\times10^3\ kg\times3.4\ \dfrac{m}{s^2}=2.72\times10^4\ N."


(b)

"v=v_0+at=0+3.4\ \dfrac{m}{s^2}\times5\ s=17\ \dfrac{m}{s}."